OpenStudy (anonymous):

A progression has a first term of 12 and a fifth term of 18. Find the sum of the first 25 terms in the progression is aritmetic.

OpenStudy (akashdeepdeb):

If th progression is an AP! Then a = 12 and a+4d = 18 So d = 6/4 = 3/2 = 1.5 So sum can be found by S = n/2(2a+(n-1)d) You have a = 12 d=1.5 n = 25 Just put it in the equation and then SOLVE!! :) Getting me? :)

OpenStudy (anonymous):

ummm im ok :) it'll take me a bit to understand could i get back to you ?

OpenStudy (anonymous):

Perfect ! i get it ! thanx a lot

OpenStudy (akashdeepdeb):

:)